OpenStudy (anonymous):

Last month, Steven had $4285.68 that he divided into two accounts. He deposited 1/3 into a savings account. He deposited the rest into his checking account. How much money did Steven deposit into his checking account?

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

he deposited 2/3 of his money into his checking account so multiply 2/3 and 4285.68

OpenStudy (anonymous):

so 1.5*4285.68= 6428.52

OpenStudy (aripotta):

not 1.5, where did you get that from? 2/3(4285.68

OpenStudy (aripotta):

)

OpenStudy (anonymous):

3 divided 2=1.5 I was trying to make it into a decimal

OpenStudy (aripotta):

2/3 is 2 divided by 3, not 3 divided by 2. and 2/3 = 0.667

OpenStudy (anonymous):

oops...So .667*4285.68= 2858.54856

OpenStudy (anonymous):

I need to round right?

OpenStudy (anonymous):

so it would be 2858.55?

OpenStudy (aripotta):

if you do it in a calculator 2 divided by 3 first, then the equal sign, and then multiply the answer with 4285.68, you should get 2857.12

OpenStudy (anonymous):

So the answer is 2857.12

OpenStudy (aripotta):

yea

OpenStudy (aripotta):

they might want that in a fraction tho

OpenStudy (aripotta):

idk

OpenStudy (anonymous):

Thank you i have just one more think you can help?

OpenStudy (aripotta):

lol thought this was the last one

OpenStudy (aripotta):

but yes, i'll help :)

OpenStudy (anonymous):

so did I